OpenStudy (anonymous):

What are the possible number of negative zeros of f(x) = 2x^7 – 2x^6 + 7x^5 – 7x^4 – 4x^3 + 4x^2 ? i got 1 not sure help?

OpenStudy (campbell_st):

take x^2 as a factor \[P(x) = x^2(2x^5 - 2x^4 + 7x^3 - 7x^2 - 4x + 4)\] x - 1 is a factor since P(1) = 1(2 - 2 + 7 - 7 - 4 + 4) P(1) = 0 which gives \[P(x) = x^2(x -1)(2x^4 + 7x^2 - 4) \] \[(2x^4 + 7x^2 - 4) = (2x^2 - 1)(x^2 +4) \] so \[P(x) = x^2(x+1)(2x^2 -1)(x^2 + 4) \]

OpenStudy (anonymous):

1 2 or 0 4, 2, or 0 7, 5, 3, or 1

OpenStudy (campbell_st):

so there are 7 possible zeros

OpenStudy (anonymous):

7,5,3, or 1

OpenStudy (anonymous):

is that the answer

OpenStudy (campbell_st):

no I just read the negative bit \[P(x) = x^2 (x -1)(2x^2 -1)(x^2 +4) \] then there are 2

OpenStudy (anonymous):

so its 2 or 0

OpenStudy (anonymous):

I was going to tell you that (x+1) in your previous post should be (x-1)

OpenStudy (anonymous):

? can you help with this question please

OpenStudy (anonymous):

Actually there is only one that is < 0 \[ x=-\sqrt{\frac 1 2} \]

OpenStudy (anonymous):

im really confused everyone is giving me a different answer

OpenStudy (campbell_st):

what about the complex from x^2 + 4

OpenStudy (anonymous):

These are not negative numbers. Negative numbers are only real numbers.

OpenStudy (anonymous):

x^2 + 4 > 0 and never zero

OpenStudy (anonymous):

for x real.

OpenStudy (anonymous):

A 1 B 2 or 0 C 4, 2, or 0 D 7, 5, 3, or 1 ????????????????????????

OpenStudy (anonymous):

A is your answer. You can take it to the bank.

OpenStudy (anonymous):

thats what i got first

OpenStudy (anonymous):

You are right. There is only one negative real root.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw